A mixture contains carbon dioxide and carbon monoxide its molecular mass is 40 gram per mole find the mole percentage of Co2 carbon dioxide

Grade:12th pass

1 Answers

Arun
5 years ago
molecular weight of mixture of CO and CO2= 40,   Let % of mole of CO be a in mixture then Average molecular weight = a × 28 + (100 – a) 44 / 100,   40 = 28a + 4400 – 44a / 100,   a = 25   Mole % of CO = 25   Mole % of CO2= 75